OpenStudy (anonymous):

Test teh following sequences for convergence or divergence. I will put teh series with teh equation box

OpenStudy (anonymous):

Give me a miin to put up teh equation

OpenStudy (phi):

what does teh mean?

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

{1/ n ln (n)} from n=2 to infinity

OpenStudy (phi):

is the ln(n) in the numerator or denominator?

OpenStudy (anonymous):

the numerator is 1 and the denominator is n*ln(n)

OpenStudy (phi):

You have to apply convergence tests.

OpenStudy (anonymous):

we just started doing series so im not too sure how to approach the problem nor what a convergence test is

OpenStudy (anonymous):

can u explain

OpenStudy (phi):

which ones are you learning?

OpenStudy (anonymous):

we only touched upon the nth term divergence test

OpenStudy (phi):

which is?

OpenStudy (anonymous):

if the limit n approaches infinity of a sub n is nto equal to 0, then \[\sum_{n=1}^{\infty}\] diverges

OpenStudy (anonymous):

Also we learned if teh limit of the nth term of {a sub n} exist, then (a sub n} converges to the limit. if it doesnt exist, then it is divergent

OpenStudy (anonymous):

Thats basically what we learned so far

OpenStudy (phi):

This series needs something more powerful. \[\sum_{1}^{\infty}a_n \text{ converges if and only if } \sum_{1}^{\infty}2^ka_{2k}\text{ converges}\]

OpenStudy (phi):

\[\sum_{k=1}^{\infty}\frac{2^k}{2^k \ln(2^k) }=\sum_{k=1}^{\infty}\frac{1}{ \ln(2^k) }=\sum_{k=1}^{\infty}\frac{1}{ k \ln(2) }=\frac{1}{\ln2}\sum_{k=1}^{\infty}\frac{1}{ k}\]

OpenStudy (anonymous):

OMG is this the snwer

OpenStudy (phi):

that last series is the harmonic series. It is also nasty to analyze (caused a lot of headaches back in the day). I think you use the same 2^k idea. But it does not converge.

OpenStudy (phi):

It does diverge

OpenStudy (anonymous):

i mean why does it diverge

OpenStudy (phi):

Do you know the integral test?

OpenStudy (anonymous):

i dont rember actually

OpenStudy (phi):

\[\sum_{k=1}^{\infty}\frac{1}{k}-->\sum_{k=1}^{\infty}\frac{2^k}{2^k}=\sum_{k=1}^{\infty}1\]

OpenStudy (phi):

That's the Cauchy test. The harmonic series diverges, so your original series diverges

OpenStudy (anonymous):

um that doenst look familiar

OpenStudy (anonymous):

First question, a very important one. You have the words "sequence" and "series" in your original question. Are these sequences or series? Their convergences are completely different.

OpenStudy (phi):

I don't know all the ways to prove this stuff. But I know the ratio test, and the root test don't help.

OpenStudy (anonymous):

Oh its a sequence

OpenStudy (phi):

You mean we do not add them up!!

OpenStudy (anonymous):

Did u add them

OpenStudy (phi):

that is what the summation sign means, add up all the terms up to infinity

OpenStudy (anonymous):

well the original question is : test the following sequences for convergence or divergence And the expresion is in brackets.

OpenStudy (phi):

In that case, the sequence converges, because when n --> infinity the nth term approaches zero

OpenStudy (phi):

Sorry to have dragged you through all the nasty math.

OpenStudy (anonymous):

OH NO ....im soo sorry taht i made u do more work than necessary. its my fault

OpenStudy (anonymous):

im soo glad u are helpingme

OpenStudy (anonymous):

So is there any work that can be shown other than that explanation that nth term approaches 0

OpenStudy (phi):

That's all you need to say

OpenStudy (anonymous):

so does teh n=2 really matter

OpenStudy (phi):

\[ \frac{1}{big number} -->0 \]

OpenStudy (anonymous):

that makes sense

OpenStudy (phi):

the ln(1)=0 , and it is not good to divide by 0, so they start with 2

OpenStudy (anonymous):

wat does the N and Ns mean

OpenStudy (anonymous):

or wat does that expression mean in words

OpenStudy (anonymous):

can u help me determine the convergence/divergence of 2 more sequences after this

OpenStudy (anonymous):

Ignore my math jargon. A sequence converges to a number L if \[\lim_{n \rightarrow \infty} a_n \rightarrow L\] For example, given the sequence \[a_n=\frac{1}{n}\] \[\lim_{n \rightarrow \infty} \frac{1}{n} \rightarrow 0; (a_n) \rightarrow 0\]

OpenStudy (anonymous):

oh so since it goes 2 0, it converges right?

OpenStudy (anonymous):

Yup. It only doesn't converge if the limit doesn't exist/diverges.

OpenStudy (anonymous):

ok cool i see. that is teh proof then. can u help me w/ 2 more???

OpenStudy (phi):

did you post it

OpenStudy (anonymous):

no i was just waitng to see if u guys dont mind helping

OpenStudy (phi):

post it,